# Rudin-Shapiro function along irreducible polynomials over finite fields

### v1

@@ 2.4. Vaughan’s identity for function fields
In this section we revise Vaughan’s identity for function fields. To do that, let us define von Mangoldt function $\Lambda$ in $\mathbb{F}_{q}[t]$ by
@@
$\mu(f)=\left\{\begin{array}[]{cl}(-1)^{k}&\text{if $f=P_{1}\dots P_{k}$ for some distinct }\\ &\text{\quad irreducible polynomials $P_{1},\dots,P_{k}$;}\\ 0&\text{otherwise.}\end{array}\right.$
@@ 3. Proof of Theorem 1
(7) ${\mathcal{R}}(\gamma)=\frac{1}{q}\sum_{\psi}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f)-\gamma),$
@@
(8) $\frac{1}{q}\sum_{f\in{\mathcal{P}}(n)}\psi_{0}(R(f)-\gamma)=\frac{\#{\mathcal{P}}(n)}{q}$
@@
(9) $\begin{split}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))&=\frac{1}{n}\sum_{f\in{\mathcal{P}}(n)}\Lambda(f)\psi(R(f))\\ &=\frac{1}{n}\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))+O(q^{n/2}).\end{split}$
Put
(10) $u=\left\lceil 3n/10\right\rceil\quad\text{and}\quad v=\left\lceil 19n/30\right\rceil.$
@@
$\Sigma_{2}=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\psi(R(hg_{1}))\overline{\psi(R(hg_{2}))}\right|.$
In order to estimate $\Sigma_{1}$, observe, that by Lemma 4, the quadratic polynomial $h\mapsto R(gh)$ has rank at least $n-\deg g-1$. Then, by Lemma 2,
(12) $\Sigma_{1}$ $\ll\sum_{|g|\leq q^{u+v}}q^{n-\deg g-(n-\deg g)/2}$
$=\sum_{|g|\leq q^{u+v}}q^{(n-\deg g)/2}$
$\ll q^{n/2}\sum_{i\leq u+v}q^{i/2}$
$\ll q^{(n+u+v)/2}\ll q^{29n/30}$
using (10).
@@
$S(g_{1},g_{1})=\sum_{|h|=q^{i}}\psi(R(g_{1}h)-R(g_{2}h)).$
with monic polynomials $g_{1},g_{2}$ of degree $n-i$. For any $g_{2}$, the number of $g_{1}$ such that the bilinear form $h\mapsto R(g_{1}h)-R(g_{2}h)$ has rank at most $n-2\deg g_{1}=2i-n$ is $O(q^{(n-i)/2})$ by Lemma 3. For those polynomials $g_{1}$, we estimate $S(g_{1},g_{2})$ trivially by $q^{i}$. Otherwise, we apply Lemma 2 to get
$S(g_{1},g_{2})\ll q^{i-(2i-n)/2}=q^{n/2}.$
Then
$\Sigma_{2}\ll\max_{v\leq i\leq n-u}\left(q^{(n-i)/2}q^{i}+q^{n-i}q^{n/2}\right)\ll q^{n-u/2}+q^{3n/2-v}.$
It yields
(13) $n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2}$ $\ll n^{5/2}\left(q^{3n/2-3u/4-v/2}+q^{7n/4-u/2-v}\right)$
$\ll n^{5/2}q^{29n/30}$
using (10). Then, it follows from (9), (11), (12) and (13) that
$\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))\ll n^{3/2}q^{29n/30}.$
Then, the result follows from (7) and (8).

### v2

@@ 4.4. Vaughan’s identity for function fields
In this section we revise Vaughan’s identity for function fields. To do that, let us define the von Mangoldt function $\Lambda$ in $\mathbb{F}_{q}[t]$ by
@@
[…]
@@ 5. Proof of Theorem 1
${\mathcal{R}}(\gamma)=\frac{1}{q}\sum_{\psi}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f)-\gamma),$
@@
$\frac{1}{q}\sum_{f\in{\mathcal{P}}(n)}\psi_{0}(R(f)-\gamma)=\frac{\#{\mathcal{P}}(n)}{q}$
@@
$\begin{split}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))&=\frac{1}{n}\sum_{f\in{\mathcal{P}}(n)}\Lambda(f)\psi(R(f))\\ &=\frac{1}{n}\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))+O(q^{n/2}).\end{split}$
@@
$\Sigma_{2}=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\psi(R(hg_{1}))\overline{\psi(R(hg_{2}))}\right|$
with some $u,v$ to be chosen later.
In order to estimate $\Sigma_{1}$, observe, that by Lemma 7, the quadratic form $h\mapsto S(gh)$ with $\deg h\leq i$ has rank at least $n-\deg g-1$. If we consider it only on monic polynomials $h$, then its rank reduced by at most one and in this case $R(gh)=S(gh)+L_{g}(h)$ by (2) with some linear polynomial $L_{g}$ in terms of the coefficients of $h$. Then, by Lemma 4,
(15) $\Sigma_{1}$ $\ll\sum_{|g|\leq q^{u+v}}q^{n-\deg g-(n-\deg g-2)/2}$
$=\sum_{|g|\leq q^{u+v}}q^{(n-\deg g)/2+1}$
$\ll q^{(n+2)/2}\sum_{i\leq u+v}q^{i/2}$
$\ll q^{(n+u+v+2)/2}.$
@@
$I(g_{1},g_{1})=\sum_{|h|=q^{i}}\psi(R(g_{1}h)-R(g_{2}h)).$
with monic polynomials $g_{1},g_{2}$ of degree nin-i. For given $g_{1}$, let
${\mathcal{B}}(g_{1})=\{g\in{\mathcal{M}}(n-i):g^{*}g=g_{1}^{*}g_{1}\}.$
If $g_{1}\not\in{\mathcal{B}}(g_{1})$, we have by Lemma 8, that the rank of the quadratic form $h\mapsto S(g_{1}h)-S(g_{2}h)$ with $\deg h\leq i$ is at least n2(ni)1=2in1n-2(n-i)-1=2i-n-1. If we consider it only on monic polynomials $h\in{\mathcal{M}}(i)$, then its rank is reduced by at most one and in this case $R(g_{1}h)-R(g_{2}h)=S(g_{1}h)-S(g_{2}h)+L_{g_{1},g_{2}}(h)$ with some linear polynomial $L_{g_{1},g_{2}}$ in the coefficients of $h$. Thus, in this case by Lemma 4 we get
$I(g_{1},g_{2})\ll q^{i-(2i-n-2)/2}=q^{n/2+1}.$
If $g_{2}\in{\mathcal{B}}(g_{1})$, we use the trivial bound
$I(g_{1},g_{2})\ll q^{i}.$
By Lemma 6 we have
${\mathcal{B}}(g_{1})\ll\min\left\{2^{n},q^{(2+\varepsilon)\frac{n}{\log_{q}n}}\right\}.$
By choosing $\varepsilon=1/2$, we get
${\mathcal{B}}(g_{1})\ll q^{n/14},$
where the implied constant is absolute. Then
(16) $\Sigma_{2}\ll\max_{v\leq i\leq n-u}\left(q^{n/14}q^{i}+q^{n-i}q^{n/2+1}\right)\ll q^{15n/14-u}+q^{3n/2-v+1}.$
Combining (14), (15) and (16), we get
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))$ $\ll nq^{(n+u+v+2)/2}$
$\quad+n^{5/2}q^{n-(u+v)/2}\left(q^{15n/28-u/2}+q^{3n/4-v/2+1/2}\right)$
$\ll nq^{(n+u+v+2)/2}$
$\quad+n^{5/2}\left(q^{43n/28-u-v/2}+q^{7n/4-u/2-v+1/2}\right).$
Choosing
$u=\frac{3}{14}n\quad\text{and}\quad v=\frac{10}{14}n,$
we get
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))$ $\ll n^{5/2}q^{27n/28+1}.$